### v2

@@ 3.3. Invisible Robot
The AR robot’s 1:1 scale with the real Franka Panda robot often occludes the surroundings. To address this, the invisible robot visualization renders the AR robot invisible and displays a cylinder at the end effector’s position. The cylinder’s color changes based on the gripper’s state: green when the gripper is open and red when the gripper is closed (see Figure 2). The color provides users with intuitive visual feedback about the gripper’s current state without obstructing their view of the environment.
@@
Figure 3. User comments about the six demonstration collection interfaces used in the formative study.
While prior work has compared various demonstration collection interfaces, there is a lack of research on demonstration collection with AR. Thus, our formative study examines the experiences and needs of end users with AR-based collection interfaces.
@@ 4.1. Method
Participants. Since EVE aims to support everyday users, we recruited 10 participants via snowball sampling. Participants were screened via a demographic questionnaire, which asked about their demographics and backgrounds, including age, gender, area of expertise, familiarity with robotics (5-point Likert scale), familiarity with demonstration collection interfaces (5-point Likert scale), and tasks they prefer the robot to help them with. We followed best practices from the HCI Guidelines for Gender Equity and Inclusivity (Scheuerman et al., 2020) to collect gender-related information. Participants ($\mu=20.7$ years old, $\sigma=1.7$ years), denoted as F01-F10, represented seven different areas of expertise and varying levels of familiarity in robotics and demonstration collection (Table A). The participants’ familiarity with robotics ranged from unfamiliar to somewhat familiar: one person was unfamiliar, five were somewhat unfamiliar, three had neutral familiarity, and one was somewhat familiar. Regarding their familiarity with the demonstration collection, six participants were unfamiliar, and four were somewhat unfamiliar.
@@
The formative study consisted of a two-part demonstration collection with six demonstration collection interfaces. The researchers began by demonstrating how to use each interface. Then, participants were given a few minutes to try each interface and ask any technical questions. Afterward, participants were given up to five minutes to try each interface by pushing, picking, and placing various cubes. Since we are investigating the user experience with various collection interfaces, we did not design the task to simulate a real-world scenario.
In part one, participants used kinesthetic teaching, teleoperation with an HTC Vive VR controller, and the AR2-D2 system. After the participants had tried data collection with the three state-of-the-art collection interfaces, we conducted a semi-structured interview about their experience, focusing on what features they liked and disliked about each interface. Although part one does not provide direct insight into an AR-based demonstration collection system, we included it to provide participants with a more comprehensive experience with robot data collection. In part two, participants performed the same task with AR-KT, path history, and invisible robot visualizations. Aside from discussing the same questions in part one, we brainstormed additional features for the AR-based collection interface. Subsequently, participants completed a survey for System Usability Scale (SUS) scores and a form to rank the motion intent communication, user enjoyment, and preference for the AR2-D2 system and the new AR visualizations. Finally, we explained to the participant how the participant’s data would be used to investigate the design for AR-based demonstration collection interfaces and answered any relevant questions.
Analysis. We analyzed two sources of data: interview transcripts and the post-task questionnaires. For the qualitative data, we used reflexive thematic coding to analyze the usability challenges faced by the participants with the current AR-based demonstration collection system. The first author, who facilitated all user study sessions, created an initial codebook by reviewing study transcripts. The entire team then collaboratively iterated on the codebook while checking for bias and coverage. With a final codebook consisting of 15 codes, the team discussed the resulting themes. While the formative study focuses on participants’ reactions to AR-based collection interfaces, we summarized and selected several participants’ quotes regarding what they liked and disliked about each collection interface (Figure 3). For SUS scores, we converted survey responses, which are on a scale of 0-40 when summed, to a range between 0-100(Brooke, 1996)111(((Q1 + Q3 + Q5 + Q7 + Q9) - 5) + (25 - (Q2 + Q4 + Q6 + Q8 + Q10))) * 2.5. We then conducted a Friedman test as an omnibus test with an appropriate number of Wilcoxon signed-rank tests corrected with Holm’s sequential Bonferroni procedure for statistical significance. See Table 1 for a summary of quantitative results from the post-task questionnaires.
@@
Usability (SUS) 66.3 (20.3) 67.3 (12.3) 70.0 (13.0) 68.3 (15.7)
Motion Intent (rank) 2.9 (1.2) 2.4 (1.0) 1.9 (0.9) 2.8 (1.3)
Enjoyment (rank) 3.5 (1.0)^ 2.4 (1.0) 2.0 (0.7)^ 2.1 (1.3)
Preference (rank) 3.5 (1.0)^ 2.7 (0.8) 1.5 (0.7)^ 2.3 (1.1)
Table 1. Mean and standard deviation of usability, motion intent communication, enjoyment, and preference for different AR visualizations. Usability is rated on a scale from 0 to 100, with higher scores indicating better usability. Motion intent communication, user enjoyment, and overall user preference are ranked from 1 to 4, with lower ranks indicating better performance. For statistical significance, one caret (^) is $p<0.05$.
@@ 4.2. Findings
A Friedman test found no statistically significant differences between the four AR visualizations in the SUS scores ($\chi^{2}(3)=2.6$, $p=0.5$) and rankings of motion intent communication ($\chi^{2}(3)=3.7$, $p=0.3$). The path history visualization significantly outperformed the baseline AR2-D2 system in user enjoyment ($p=0.0059$) and preference ($p=0.0059$). Notably, the path history visualization achieved the highest median scores and rankings across all measured variables: SUS ($\mu=70.0$, $\sigma=13.0$), motion intent communication ($\mu=1.9$, $\sigma=0.9$), user enjoyment ($\mu=2.0$, $\sigma=0.7$), and preference ($\mu=1.5$, $\sigma=0.7$). These findings suggest that participants found the path history visualization the most usable, enjoyable, and preferable among the four tested AR visualizations, even though it did not significantly outperform the others in all aspects. Interestingly, these results contradicted our hypotheses. Given the benefits of direct manipulation and real-time visual feedback, we had expected the AR-KT visualization to achieve the highest scores on system usability and motion intent communication compared to other visualizations. The contradiction may be due to the AR robot’s inability to accurately track the hand when it moves quickly. We also anticipated that the invisible robot visualization would receive the lowest scores on system usability and motion intent communication due to the lack of joint movement information. The contradiction could be attributed to the approximate straight-line motion of the AR robot, which allows users to easily predict the movement of the joints.
@@
    Ambiguous spatial position of the hand and the robot: Four participants (F03, F04, F08, F10) expressed difficulty interpreting the 3D environment via a 2D iPad screen. The lack of depth perception made it challenging for users to accurately gauge the relative positions of their hands and the AR robot. This ambiguity led to inaccuracies in the collected demonstrations. F03 also complained about the hand ”passing through” the robot, resulting in an unnatural collection.
@@
    Imprecise trajectory visualization: The path history visualization displays a straight green line between the robot’s end effector coordinate and the user’s hand coordinate. Although three participants (F05, F06, F10) have asserted that the trajectory visualization aids their understanding of the robot’s motion intent, three other participants (F02, F04, F07) have proposed to include a more accurate trajectory.
@@
    Absence of feedback regarding the collection efficacy: Five participants (F01, F04, F08, F09, F10) complained about the lack of collision detection, which prevented them from verifying whether the collected trajectory was feasible. F01 claimed that while AR is ”useful and fun,” it does not make sense for the AR robot to pass through all the obstacles.
@@
    Inconsistent hand tracking: Inconsistencies in hand tracking emerged as a significant concern among participants. Five individuals (F01, F02, F03, F06, F08) highlighted the presence of inaccuracies and latency in the hand tracking system employed in the AR environment. Furthermore, F02 noted that the gray line renderer, utilized in the AR2-D2 condition to visualize the tracked coordinates of the user’s fingers, proved to be visually distracting.
5. EVE Prototype 2
Figure 4. Overview of evaluation user study and prototype 2 of EVE. Left: the study involved three common tabletop tasks: toggling a switch, sorting food, and sweeping the table. These tasks were utilized to evaluate the user study with various baseline interfaces. To ensure a fair comparison of performance across these interfaces, the positions of all target objects and desired actions were fixed. Right: this section presents a sequence of frames from a sampled rollout performed by users during the evaluation. It illustrates how participants utilized the visualization and features of EVE prototype 2 to successfully complete the required demonstrations for the given task.
Informed by formative study findings and our own experiences using EVE, we created a second EVE prototype with seven additional system changes to address the usability challenges. We describe these advancements below:
@@ 5.1. Hand Position Projection
To communicate the position of the tracked hand pose in 3D space clearly (C1), a green line is generated on the surface. We have tried to attach the green line directly to the hand but found increased visual distraction and no improvement in position interpretation. To enhance motion intent communication and reduce potential collisions, the length and width of the projected line were adjusted to match the dimensions of the robot’s end effector. This improved representation of the hand position also helps prevent users from specifying coordinates that are clearly beyond the robot’s joint limits (C2).
@@ 5.2. Joint Constraints Signifier
To address user concerns about joint constraints in AR2-D2 and EVE prototype 1 (C2), we followed the joint space limits provided in the Franka Emika Panda robot specifications to define a range of reachable coordinates based on the instantiation point of the AR robot. When the user’s hand is positioned outside the reachable range, all the app interface buttons turn red (otherwise white) and are deactivated, preventing the user from changing the robot’s state.
@@ 5.3. Realistic Trajectory
Three participants (F02, F04, F07) indicated that the inaccurate, straight-line trajectory visualization did not effectively communicate the robot’s motion intent ((C3)). In response, the system was modified to display the precise, curved trajectory of the end effector. Moreover, Path History is now the default visualization for EVE, as it demonstrated superior performance compared to other AR visualizations in the formative user study.
@@ 5.4. Trajectory Replay
The absence of feedback regarding collision detection and path feasibility significantly hinders the system’s usability (C4). However, planning a trajectory for the AR robot to avoid physical obstacles often requires constructing a depth map of the world, which can be computationally expensive for the iPad. After brainstorming potential solutions with the formative study participants, we added a replay feature demonstrating the collected trajectory and changes in the gripper state using a partially transparent robot. This feature makes it easier to determine whether the robot collided with a physical object during collection. This implementation was also inspired by some participants’ preferences (F02, F03) for the replay feature available in kinesthetic teaching. To integrate trajectory visualization and enable replay for AR-KT, we experimented with various numbers of waypoints and decided to use 20 waypoints for the trajectory collected with AR-KT.
@@ 5.5. Dynamic Camera View
To address the issue of the AR robot occupying a significant portion of the screen (C5), a formative study participant (F05) proposed enabling the user to move the camera to inspect the 3D world. However, collecting demonstrations from multiple camera angles is not ideal because it is crucial to ensure that the hand-eye calibration between the robot action and the 3D voxels is aligned for a successful transfer to real-world deployment. To address this, we save the initial camera extrinsic values after the robot has been instantiated and make the buttons disappear when the current extrinsic values differ from the saved values by more than a small threshold. This approach allows the user to freely inspect the scene from multiple angles while ensuring the demonstration is collected from a consistent camera angle. This feature also addresses the lack of information about joint constraints (C2) by allowing users to visually inspect the robot’s state and avoid specifying potentially unreachable coordinates.
@@ 5.6. Accurate Gripper Control
In AR2-D2, a gray line is rendered between the thumb and index fingers to enable users to toggle the gripper opening by adjusting the distance between the fingers. Due to limitations in the iPad camera and ARKit’s HandTrackingProvider, hand tracking is inconsistent, and toggling the gripper state is extremely unpredictable (C6), as reported by five participants from the formative study (F01, F02, F03, F06, F08). To maintain the system’s ubiquity, we did not modify the hardware. Instead, we addressed the issue of the flashing gray line renderer by removing it and adding a button to toggle the gripper opening accurately. This change ensures a more reliable and predictable user experience when controlling the gripper state.
@@ 5.7. Robot Instantiation Reselection
Our experience using the system and the feedback from one participant (F09) suggest that the inability to change the instantiation point of the robot interferes with the system’s usability, as we need to quit and restart the app every time we instantiate the robot at an undesired place. We changed the robot instantiation process into two steps to address this issue. First, a green cylinder appears on the screen, signifying where the robot will spawn. The user is then prompted with two choices: 1) instantiate the robot at the position indicated by the cylinder, or 2) reselect a position for the green cylinder. The ability to reselect the instantiation point for the AR robot greatly increases the system’s usability.
6. evaluation user study
We conducted a user study to evaluate the effectiveness of EVE compared to baseline interfaces for three common tabletop tasks. These tasks were designed based on the results of an extensive survey from BEHAVIOR-1K (Li et al., 2024) asking participants, ”What do you want robots to do for you?” We included obstacles in all tasks to simulate real-world environments where tables are often cluttered. The three tasks were as follows:
  • Toggle switch: Participants guided the robot to push a switch and turn on a light bulb.
  • Sort food: Participants were required to grab a melon and place it inside a yellow bowl, and then grab a steak and put it into a basket. This task simulated the sorting of different items into specific locations.
  • Sweep table: Participants used a sponge to sweep beans to a location marked by tape. This task simulated the cleaning of debris from a table.

### v3

@@ 3.4. Invisible robot
The AR robot’s 1:1 scale with the real Franka Panda robot often occludes the surroundings. To address this, invisible robot renders the AR robot invisible and displays a cylinder to represent the position and orientation of the end effector. The cylinder’s color changes based on the gripper’s state: green when the gripper is open and red when the gripper is closed (Figure 2). The color provides users with intuitive visual feedback about the gripper’s current state without obstructing their view of the environment.
@@ 4. Formative study
Table 1. User comments about the six demonstration collection methods used in the formative study.
[…]
User feedback on six demonstration collection interfaces in the formative study. The table includes an image of each interface alongside positive and negative comments from participants.
To study the utility of our initial prototype, we designed a formative study to 1) identify the challenges users encounter when collecting demonstrations using common collection interfaces and 2) evaluate the effectiveness of the new AR visualizations in aiding users with demonstration collection.
@@ 4.1. Method
Participants. Since EVE aims to support everyday users, we recruited $10$ participants via snowball sampling. Participants were screened via a demographic questionnaire, which asked about their demographics and backgrounds, including age, gender, area of expertise, familiarity with robotics (5-point Likert scale), familiarity with demonstration collection interfaces (5-point Likert scale), and tasks they prefer the robot to help them with. We followed best practices from the HCI Guidelines for Gender Equity and Inclusivity (Scheuerman et al., 2020) to collect gender-related information. Participants ($\mu=20.7$ years old, $\sigma=1.7$ years), denoted as F01-F10, represented seven different areas of expertise and varying levels of familiarity in robotics and demonstration collection (Table A). The participants’ familiarity with robotics ranged from unfamiliar to somewhat familiar: one person was unfamiliar, five were somewhat unfamiliar, three had neutral familiarity, and one was somewhat familiar. Regarding their familiarity with the demonstration collection, six participants were unfamiliar, and four were somewhat unfamiliar. Eight participants identified as male, while two participants identified as female.
@@
The researchers began by demonstrating how to use each interface. Then, participants were given a few minutes to try each interface and ask any technical questions. Afterward, participants were given up to five minutes to try each interface by pushing, picking, and placing various cubes.
The formative study consisted of two parts. In part one, participants used kinesthetic teaching, teleoperation with an HTC Vive VR controller, and the AR2-D2 system. After the participants had tried data collection with the three state-of-the-art collection interfaces, we conducted a semi-structured interview about their experience, focusing on what features they liked and disliked about each interface. Although part one does not provide direct insight into an AR-based demonstration collection system, we included it to provide participants with a more comprehensive experience with robot data collection.
In part two, participants performed the same task with AR-KT, path history, and invisible robot visualizations. Aside from discussing the same questions in part one, we brainstormed additional features for the AR-based collection interface. Subsequently, participants completed a survey for System Usability Scale (SUS) scores and a form ranking the motion intent communication, user enjoyment, and preference for the AR2-D2 system and the new AR visualizations. Finally, we explained to the participant how the participant’s data would be used to investigate the design for AR-based demonstration collection interfaces and answered any relevant questions.
Analysis. We analyzed two sources of data: interview transcripts and the post-task questionnaires. For the qualitative data, we used reflexive thematic coding to analyze the usability challenges faced by the participants with the current AR-based demonstration collection system. The first author, who facilitated all user study sessions, created an initial codebook by reviewing study transcripts. The entire team then collaboratively iterated on the codebook while checking for bias and coverage. With a final codebook consisting of $15$ codes, the team discussed the resulting themes. We selected and summarized several participants’ quotes regarding what they liked and disliked about each collection interface (Figure 1). For SUS scores, we converted survey responses, which are on a scale of 0-40 when summed, to a range between $0-100$(Brooke, 1996)111$(((Q1+Q3+Q5+Q7+Q9)-5)+(25-(Q2+Q4+Q6+Q8+Q10)))*2.5$. We calculated the mean and the standard deviation of the ranking forms. For all the scores, we conducted a Friedman test as an omnibus test with an appropriate number of Wilcoxon signed-rank tests corrected with Holm’s sequential Bonferroni procedure for statistical significance. See Table 2 for a summary of quantitative results from the post-task questionnaires.
@@
[…]
Table 2. Path history outperformed other AR-based collection methods in usability, motion intent communication, user enjoyment, and user preference. The mean and standard deviation of all measurements are shown. Usability is rated on a scale from 0 to 100, with higher scores indicating better usability. Motion intent communication, user enjoyment, and user preference are ranked from 1 to 4, with lower ranks indicating better performance. Statistical significance is indicated by one caret (^) for $p<0.05$.
[…]
Mean and standard deviation of usability, motion intent, enjoyment, and preference of AR2-D2, AR Kinesthetic Teaching, Path History, and Invisible Robot.
@@ 4.2. Findings
The path history visualization significantly outperformed the baseline AR2-D2 system in user enjoyment ($p=0.0059$) and user preference ($p=0.0059$). Notably, the path history visualization achieved the highest median scores and rankings across all measured variables: SUS ($\mu=70.0$, $\sigma=13.0$), motion intent communication ($\mu=1.9$, $\sigma=0.9$), user enjoyment ($\mu=2.0$, $\sigma=0.7$), and preference ($\mu=1.5$, $\sigma=0.7$). A Friedman test found no statistically significant differences between the four AR visualizations in the SUS scores ($\chi^{2}(3)=2.6$, $p=0.5$) and rankings of motion intent communication ($\chi^{2}(3)=3.7$, $p=0.3$). These findings suggest that participants found the path history visualization the most usable, enjoyable, and preferable among the four tested AR visualizations, even though it did not significantly outperform the others in all aspects.
Interestingly, these results contradicted our hypotheses. Given the benefits of direct manipulation and real-time visual feedback, we had expected the AR-KT visualization to achieve the highest scores on system usability and motion intent communication compared to other visualizations. The contradiction may be due to the AR robot’s inability to accurately track the hand when it moves quickly due to the computational limit of the device and the inaccuracy of the device’s LiDAR camera. We also anticipated that the invisible robot visualization would receive the lowest scores on system usability and motion intent communication due to the lack of joint information. The contradiction could be attributed to the approximate straight-line motion of the AR robot, which allows users to easily predict the movement of the joints.
@@
    Ambiguous spatial position of the hand and the robot: Four participants (F03, F04, F08, F10) expressed difficulty interpreting the 3D environment via a 2D iPad screen. The lack of depth perception made it challenging for users to accurately gauge the relative positions of their hands and the AR robot. This ambiguity led to inaccuracies in the collected demonstrations. F03 also complained about the hand passing through the robot, resulting in an unnatural collection.
@@
    Imprecise trajectory visualization: The path history visualization displays a straight green line between the robot’s end effector coordinate and the user’s hand coordinate. Although three participants (F05, F06, F10) stated that the trajectory visualization aids their understanding of the robot’s motion intent, three other participants (F02, F04, F07) have proposed to include a more accurate trajectory.
@@
    Absence of feedback regarding the collection efficacy: Five participants (F01, F04, F08, F09, F10) complained about the lack of collision detection, which prevented them from verifying whether the collected trajectory was feasible. F01 claimed that while AR is “useful and fun,” it does not make sense for the AR robot to pass through all the obstacles.
@@
    Inconsistent hand tracking: Inconsistencies in hand tracking emerged as a significant concern among participants. Five individuals (F01, F02, F03, F06, F08) highlighted the presence of inaccuracies and latency in the hand tracking system employed in the AR environment. Furthermore, F02 noted that the gray line renderer, utilized in AR2-D2 to visualize the tracked coordinates of the user’s fingers, is visually distracting.
5. EVE final prototype
Figure 3. Overview of the evaluation user study and prototype 2 of EVE. Left: The study included three common tabletop tasks: toggling a switch, sorting food, and sweeping the table. Right: An example of data collection for the food sorting task using EVE final prototype.
[…]
The setup of the three tabletop tasks used in the evaluation study (left). An example collection of the sort food task using EVE prototype 2 (right).
Informed by formative study findings and our own experiences using EVE, we upgraded EVE’s prototype with seven additional system changes to address the usability challenges. We describe these advancements below:
@@ 5.1. Hand position projection
The 2D display of the smartphone or tablet makes position perception of the hand and the robot in 3D space challenging (C1). Prior work addresses this problem with a virtual cast shadow of a flying drone (Chen et al., 2021). We imitated this approach by representing the shadow of the AR robot’s end effector with a green line on the surface of the table. To further enhance motion intent communication, the length and width of the projected line were adjusted to match the dimensions of the robot’s end effector. This feature also aids users become familiar with the robot’s morphology and estimate what positions would be unreachable (C2).
@@ 5.2. Joint constraints signifier
To address user concerns about joint constraints (C2), we adhered to the joint space limits specified in the Franka Emika Panda robot documentation (GmbH, 2024). This allowed us to define a range of reachable coordinates based on the AR robot’s initial position. When the user’s hand is outside this reachable range, all buttons turn red (otherwise white) and are deactivated, preventing any changes to the robot’s state.
@@ 5.3. Realistic trajectory
Three participants (F02, F04, F07) indicated that the straight-line trajectory visualization did not effectively communicate the robot’s motion intent (C3). In response, the system was modified to display the precise, curved trajectory of the end effector. Moreover, path history is now the default visualization for EVE, as it demonstrated superior performance compared to other AR visualizations in the formative user study. We disabled trajectory visualization during AR-KT because the user is controlling the robot in real-time.
@@ 5.4. Trajectory replay
The lack of feedback regarding collision detection and path feasibility significantly hinders the system’s usability (C4). Detecting collisions of the AR robot with the physical objects or planning a trajectory for the AR robot to avoid physical obstacles typically requires constructing a depth map of the environment, which can be computationally expensive for the iPad. After brainstorming potential solutions with the formative study participants, we added a replay feature that demonstrates the collected trajectory and changes in the gripper state using a partially transparent robot. This feature makes it easier to determine whether the robot collided with a physical object during collection. Additionally, this implementation was inspired by some participants’ preferences (F02, F03) for the replay feature available in kinesthetic teaching. The system linearly interpolates 20 waypoints for the trajectory collected with AR-KT, allowing users to modify waypoints and replay the trajectory.
@@ 5.5. Dynamic camera view
To address the issue of the AR robot occupying a significant portion of the screen (C5), users need to view the robot from different angles. Prior work enhanced situational awareness using an interactive third-person perspective from a second, spatially coupled drone (Temma et al., 2019). After discussing with a formative study participant (F05), we decided to allow users to move the device to inspect the 3D world, as setting up additional smartphones or tablets is costly and inconvenient. However, collecting demonstrations from multiple camera angles is not ideal because it is crucial to maintain hand-eye calibration between the robot’s actions and the 3D voxels for successful real-world deployment. Different camera angles might provide conflicting or misleading information about distances and spatial relationships, which can disorient the robot’s understanding and actions. To address this, we save the initial camera extrinsic values after the robot is instantiated and make the buttons disappear when the current extrinsic values deviate from the saved values by more than a small threshold. This approach enables users to freely inspect the scene from multiple angles while ensuring the demonstration is collected from a consistent camera angle. This feature also helps users become familiar with the robot’s morphology and estimate which positions might be unreachable (C2).
@@ 5.6. Accurate gripper control
In AR2-D2, a gray line is rendered between the thumb and index fingers, allowing users to toggle the gripper opening by adjusting the distance between their fingers. However, due to limitations in the iPad camera and ARKit’s HandTrackingProvider, hand tracking is inconsistent, making the toggling of the gripper state unpredictable (C6), as reported by five participants from the formative study (F01, F02, F03, F06, F08). To maintain the system’s ubiquity, we opted not to modify the hardware. Instead, we resolved the issue of the flashing gray line renderer by removing it and incorporating a button to accurately toggle the gripper opening. This adjustment ensures a more reliable and predictable user experience when controlling the gripper state.
@@ 5.7. Robot instantiation reselection
Our experience with the system, along with feedback from one participant (F09), suggests that the inability to change the robot’s instantiation point hinders usability, as it requires restarting the application if the robot is placed in an undesired location. To address this issue, we revised the robot instantiation process into two steps. First, a green cylinder appears on the screen to indicate where the robot will spawn. The user is then given two choices: instantiate the robot at the position indicated by the cylinder or reselect a position for the green cylinder. This ability to reselect the instantiation point for the AR robot significantly enhances the system’s usability.
6. Evaluation
We conducted a user study to evaluate the effectiveness of EVE compared to baseline interfaces for three common tabletop tasks. These tasks were designed based on the survey from BEHAVIOR-1K (Li et al., 2024) asking participants, “What do you want robots to do for you?” We included obstacles in all tasks to simulate real-world environments where tables are often cluttered. The three tasks were as follows:
  • Toggle switch: Participants guided the robot to push a switch, turning on a light bulb. This task is more challenging than the button press task from the AR2- D2 paper, as it requires millimeter-level precision for the robot’s end-effector to touch a specific point on the switch.
  • Sort food: Participants controlled the robot to grab a melon and place it inside a yellow bowl, then pick up a steak and put it into a basket. This task simulated sorting different items into specific locations.
  • Sweep table: Participants directed the robot to sweep beans to a location marked by tape with a sponge. This task simulated cleaning debris from a table.